# Future proofing network encryption technology (and securing critical infrastructure data) with continuous-variable quantum key distribution

### v1

@@ III.1 Dynamic polarization optimization
Controlling the SOP of the modulated signal so that it matches with the SOP of the LLO yields a stronger RF heterodyne output. The objective function to maximize was the overall power $R$ of the RF heterodyne signal, the spectrum of which is shown in Fig. 2(c) by the red trace (with the background noise contribution shown by the blue trace). We remark here that even after channel attenuation, the SNR was reasonably large, simply due to the various multiplexed ancillary signals. And finally, also note that the SOP of the LLO was fixed by virtue of using only PMF components, namely the Rx laser, LO switch, and 50:50 fiber splitter; see Fig. 1.
The optimization was executed in a feedback loop, iteratively seeking to increase the power $R_{n}$ and settle around a maximum after some $n=N$ rounds; typically $N\leq 35$ sufficed for our system. At the end of iteration $n$, ‘optimized’ digital inputs $\vec{v}_{n+1}$ for the multi-channel DAC were evaluated using digital inputs $\vec{v}_{n}$ and a gradient estimate $\vec{g}_{n}$ as $\vec{v}_{n+1}=\vec{v}_{n}-b_{n}\vec{g}_{n}$. Here, $b_{n}$ denotes a pre-calculated scalar gain coefficient, and the vector sign is to convey that the operation was on 4 channels simultaneously.
Table 2: Polarization optimization task log from one of several experiments performed at DTU Lyngby. The components of the vector $\vec{v}_{n}$ in columns $2$-$5$ are the digital inputs of four 16-bit DACs. A background noise power $R_{0}=33.831$ is already corrected for in the last column. All DACs are initialized in the beginning to $v_{1}[j]=2^{15}-1=32767$, $j\in[1,4]$, to have the maximum operating range.
$n$
$v_{n}[0]$
$v_{n}[1]$
$v_{n}[2]$
$v_{n}[3]$
$R_{n}$
1
32767
32767
32767
32767
92.863
2
38091
27443
38091
27443
65.792
3
39488
26046
36694
28840
69.726
4
39032
26502
36238
28384
72.070
5
38189
27345
37081
29227
78.407
6
36971
26127
35863
30445
90.376
7
37044
26054
35936
30518
93.489
8
36695
26403
36285
30867
98.593
9
36445
26153
36535
30617
98.523
10
36228
25936
36318
30834
100.421
11
35804
26360
35894
31258
103.931
12
35612
26552
36086
31066
104.661
13
34746
27418
35220
31932
118.385
14
35050
27114
34916
32236
113.494
15
34968
27032
34998
32318
113.732
16
34957
27021
34987
32307
112.943
17
34861
26925
35083
32403
114.292
18
34588
27198
34810
32676
115.751
19
34388
27398
35010
32476
116.653
20
34248
27538
35150
32616
118.133
21
33986
27800
35412
32878
118.465
22
34080
27706
35506
32784
118.291
23
34111
27737
35537
32815
118.177
24
33977
27603
35671
32949
118.283
25
34023
27649
35717
32995
117.585
26
34155
27781
35585
33127
118.492
27
34154
27780
35586
33126
117.902
28
34120
27814
35620
33160
119.007
29
34138
27832
35638
33142
118.035
30
34085
27779
35691
33195
118.137
31
34077
27787
35699
33203
117.726
32
34084
27794
35706
33210
118.448
33
34026
27736
35764
33268
117.808
34
33979
27783
35717
33315
117.717
35
33971
27791
35725
33323
117.888
The calculation of $\vec{g}_{n}$ during the iteration involved applying successive positive and negative perturbations $\vec{v}_{n}+p_{n}\vec{d}_{n}$ and $\vec{v}_{n}-p_{n}\vec{d}_{n}$, respectively, and measuring the corresponding powers $R^{+}_{n}$ and $R^{-}_{n}$ to yield
$\vec{g}_{n}=\left(\frac{R^{-}_{n}-R^{+}_{n}}{2p_{n}}\right)\vec{d}_{n}.$ (1)
In the above, $p_{n}$ denotes a pre-calculated scalar perturbation coefficient, while $\vec{d}_{n}$ is a vector with elements $1$ or $-1$ drawn from a Bernoulli distribution, with both elements being equally likely.
Table 2 shows the log from a typical run of the polarization optimization task. The four columns in the center are the components of $\vec{v}_{n}$, while the corresponding evolution of powers $R_{n}$ is presented in the last column. One can observe that $R_{n}$ initially goes down but quickly recovers and settles to a maximum after $n\sim 20$. The corresponding size of the perturbation kicks ($|\vec{v}_{n+1}-\vec{v}_{n}|$) follows the same behaviour. Note that at the beginning of the optimization task, we measure the background noise power $R_{0}$ (due to vacuum + electronic noise) by making use of the Sig switch, and subtract it from the total power in the RF heterodyne measurement.
The stochastic nature of SPSA is embedded in the fact that $\vec{d}_{n}$ is chosen randomly in every iteration. The choice of scalars $b_{n}$ and $p_{n}$ is critical to the performance of SPSA [31]. Through empirical tests on L1, i.e., the campus loop at DTU (see Table 1), we found that $p_{n}=p_{0}n^{-0.5}$ and $b_{n}=b_{0}n^{-0.75}$ with $p_{0}$ and $b_{0}$ in the neighbourhood of 350 and 200000, respectively, yielded a relatively fast and stable convergence, such as the one shown in Table 2. We note that there can be other combinations $p_{0}$ and $b_{0}$ that can yield a similar performance.
On links L2-L4, we faced more challenges in determining appropriate $p_{n}$ and $b_{n}$ due to which the eventual performance of the algorithm was somewhat worse. This was however not unexpected as compared to underground fibers, the rate of change of the SOP of a lightwave in OPGWs can be much higher, as they are exposed to not only larger and faster temperature variations but also the natural forces of wind and lightning [32]. With long-term testing (over few weeks or even months), we expect to be able to characterize these links more properly.
III.2 Protocol runs: Quantum phase
State machines running in a loop independently on both CVQKD-TX and CVQKD-RX used the quantum channel for transmission and reception, respectively, of the information carrying signal at 1550 nm, spectral representations of which are shown in Fig. 2(b) and (by the red trace) in Fig. 2(c). Simultaneously, the network encryptors described in subsection II.7 were set to operate continuously using secret keys generated from prior runs of the CVQKD system, implying that the quantum channel was also inundated with the 1300 nm signal.
At the RX, the task of polarization optimization, already described in subsection III.1, was followed by the ADC capturing samples of the modulated signal, using a level trigger set adaptively on the ADC channel itself that acquires the data. For this, in every round, the TX played out sequentially a total of 1000 frames, each constructed with $0.2$M unique Gaussian symbols using the DSP chain shown in Fig. 2(a). This burst from the TX lasted just around 3 mins, after which, it was left in an idle state until the start of the next round, which had a duration of 10 minutes.
While the RX was programmed to capture and store—as a file on disk—a maximum of 100 of these data frames, in practice, the yield was typically less due to timeouts (set-trigger-level too high) or due to capturing of noise between successive TX frames (set-trigger-level too low). After acquiring the data with modulation, a total of 2×2$\times$2 × 2G ADC samples for performing the task of vacuum noise calibration, described in section II.5, were captured. The switches routing the Sig and LO away from the detector for this purpose were then reverted to their original positions, and CVQKD-RX was then also left in an idle state until the next round. The duty cycle at the RX was typically between $70-80\%$.
Once the captured data files from all 6 rounds were available, DSP routines based on a machine learning framework [13] operating at DPI-RX kicked in. Using the various ancillary signals, the digital synchronization procedure was performed. In particular, the pilot tones assisted in clock and carrier recovery, while the QPSK band helped in optimal sampling [13, 15]. The latter relies on the fact that the QPSK data contained a sequence of symbols known to both TX and RX. After extracting some frame IDs, the demodulation of the QPSK band was completed. With all the frequency, phase, and temporal information obtained, the Gaussian quantum symbols could also be optimally extracted.
Figure 3 shows various results from experiments conducted on L2–L4 at Odense.
Figure 3: (Color online) Various results from the parameter estimation and calibration procedures in the tests carried out at the 3 links in Odense (refer Table 1 for more info). (a) Number of quantum states that were successfully prepared by TX as well as measured by RX. (b) Strength of the pilot tone which aids in the clock and carrier recovery. (c) Outcome of the calibration procedure performed after every round of modulation and acquisition. (d) and (e) Noise and loss attributed to Eve. (f) Noise assumed to be beyond Eve’s control. (g) Projected secret key rate (SKR) assuming the DSP and information reconciliation had been performed in real time. A reconciliation efficiency $\beta=95\%$ and FER = 0.5 was uniformly assumed in the calculations, and in the 3 cases marked by a hexagon, circle, and a triangle, it was actually also achieved (further details in the main text). Also, in all the non-marked cases, PE was performed using all the symbols of Alice and Bob.
@@ III.2 Protocol runs: Quantum phase
III.3 Protocol runs: Classical phase
@@
With the quantum stage of the protocol finished, the classical data processing steps of IR, PE, and PA were also performed offline. The IR was based on a multi-dimensional reverse reconciliation scheme and a bank of multi-edge-type low-density-parity-check codes were used [33]. Due to a limited number of code rates, high-performance IR was actually possible only in some selected rounds, three of which are indicated by the triangle, circle and hexagon shapes around the markers in Fig. 3(g). In these cases, we obtained a reconciliation efficiency $\beta\approx 95\%$ at a frame error rate (FER) in the vicinity of 0.5. In subsection IV.2, we further elaborate on this matter.
@@
Alice performed PE using her filtered symbols and the recovered quantum symbols of Bob, which, together with the variables such as the electronic and vacuum noise variances from various calibrations in RX, were shared by Bob via the authenticated channel. After calculating the covariance matrix, the total noise $\xi$ and transmittance $\eta$ were evaluated. These could be partitioned into trusted and untrusted components [14] using the vacuum noise normalization values shown in Fig. 3(c), and the optical transmittance of the receiver (including CWDM-RX) $\eta_{t}=0.33$. Figures 3(d) and (e) show the untrusted components $\eta_{u}=\eta/\eta_{t}$ of the transmittance (converted to loss in dB) and noise $\xi_{u}=\xi-\xi_{t}$ respectively, from the 6 rounds, while the trusted noise values $\xi_{t}$ themselves are shown in Fig. 3(f). As is often done in CVQKD, the noise is conveyed as whatever is in excess of the shot/vacuum noise in a relative sense, i.e., in shot noise unit (SNU). Together, these quantities aid in the calculation of $\chi_{BE}$, i.e., Eve’s Holevo information.
The results from PE enable us to estimate the SKR, given by
$\text{SKR}=B(1-\text{FER})\left(\beta I_{AB}-\chi_{BE}\right).$ (2)
The mutual information $I_{AB}$ between Alice and Bob is also calculated using the covariance matrix.
Figure 3(g) shows the SKR values for the experiments performed on L2-L4 at Odense. We emphasize that such rates can be achieved in practice under the assumptions (justified below in Section IV) that (1) the DSP, IR, and PA can be performed in real time, and (2) it is always possible to obtain $\beta=95\%$, FER $\approx$ 0.5. Based on that, the average SKR obtained from the experiments on L1 at Lyngby (not shown in Fig. 3) would be around 1.1 Mbps.
From a security perspective, the generated keys here are secure in the asymptotic regime, and assuming collective attacks by Eve [34, 22]. We remark that a composable key analysis [14] would likely not have yielded any positive SKR due to the not-too-large number of quantum states obtained in the experiments here as well as the relatively worse trusted loss ($\eta_{t}=0.33$ for qTReX versus $\eta_{t}=0.68$ in Ref. [14]).
IV Discussion
In this section, we discuss some aspects of our implementation as well as the obtained results in more detail.
IV.1 Loss: Untrusted vs Physical
@@
The channel loss values graphed in Fig. 3(e) over the 6 rounds show a considerable fluctuation, especially for OPGW×$\times 2$ and OPGW×$\times 4$, or links L2 and L4, respectively. Furthermore, comparing with the physical loss values presented in Table 1, a difference of at least 0.80.80.8\,0.8dB is evident. The latter can partly be explained due to the fixed loss coming from two extra fiber adapters that were used for connecting to the loopbacks in L2-L4; since they were absent during the OTDR measurements, their combined loss is not included in the entries of Table 1.
@@
The fluctuation on the other hand stems from the subpar performance of the SOP optimization in case of OPGWs, which was explained at the end of subsection III.1. A poor SOP reduces the mode-matching efficiency of the LLO with the optical signal (containing both the quantum data and the ancillary signals; see subsection II.5) entering CVQKD-RX, which can make the execution of the overall DSP chain at RX inefficient as well [15]. To elaborate, the SNR of the pilot tone plays an important role in ensuring a proper functioning of the machine learning framework, and when this SNR goes down, the performance of the carrier and clock recovery also worsens, causing the PE evaluated noise and/or loss to rise. This can be observed through the correlated trend shown by the green/blue markers for L2/L4 in Figs. 3(b) and 3(e). In the worst case, the machine learning framework can even completely fail during the state measurement and no symbols can be decoded; this was for instance the outcome in round 2 for link L4.
While the SOP optimization parameters can be chosen more carefully after a longer and more thorough characterization of the aerial fibers, another alternative is to use a so-called polarization hybrid [35, 36].
IV.2 Practical and real-time data processing

### v2

@@ 3.1 Dynamic polarization optimization
Controlling the SOP of the modulated signal so that it matches with the SOP of the LLO yields a stronger RF heterodyne output. The objective function to maximize was the overall power $R$ of the RF heterodyne signal, the spectrum of which is shown in Fig. 2(c) by the red trace (with the background noise contribution shown by the blue trace). We remark here that even after channel attenuation, the total SNR was reasonably large, simply due to the various multiplexed ancillary signals. And finally, also note that the SOP of the LLO was fixed by virtue of using only PMF components, namely the Rx laser, LO switch, and 50:50 fiber splitter; see Fig. 1.
The optimization was executed in a feedback loop, aiming to increase $R$ and settle around a maximum after some $N$ iterations; typically $N\sim 30$ sufficed for our system. In the supplementary section, we provide further details of how the SPSA method was exactly implemented.
3.2 Protocol runs: Quantum layer
State machines running in a loop independently on both CVQKD-TX and CVQKD-RX used the quantum channel for transmission and reception, respectively, of the information carrying signal at 1550 nm, spectral representations of which are shown in Fig. 2(b) and (by the red trace) in Fig. 2(c). Simultaneously, the network encryptors described in subsection 2.8 were set to operate continuously, using secret keys generated from prior runs of the CVQKD system, implying that the quantum channel was also inundated with the 1300 nm signal.
At the RX, the task of polarization optimization, already described in subsection 3.1 and the supplementary section, was followed by the ADC capturing samples of the modulated signal, using a level trigger set adaptively on the ADC channel itself that acquires the data. For this, the TX played out sequentially in every round a total of 1000 blocks, each containing 20 data frames, constructed using the DSP chain shown in Fig. 2(a) with $20\times 10^{4}$ unique Gaussian symbols utilized per block. This burst from the TX lasted just around 3 mins, after which, it was left in an idle state until the start of the next round, which had a duration of 10 minutes.
While the RX was programmed to capture and store—as a file on disk—a maximum of 100 of these data blocks, in practice, the yield was typically less due to timeouts (set-trigger-level too high) or due to capturing of noise between successive TX blocks (set-trigger-level too low). After acquiring the data with modulation, a total of $2\times 2=$ 4G ADC samples for performing the task of vacuum noise calibration, described in section 2.6, were captured. The switches routing the Sig and LO away from the detector for this purpose were then reverted to their original positions, and CVQKD-RX was then also left in an idle state until the next round. The duty cycle at the RX was typically between $70-80\%$.
Once the data files from all 6 rounds were available, DSP routines based on a machine learning framework [14] operating at DPI-RX kicked in. Using the ancillary signals, the digital synchronization procedure was performed. In particular, the pilot tones assisted in clock and carrier recovery, while the QPSK signal helped in optimal sampling [14, 17] by making use of a sequence of symbols known to both TX and RX. After extracting the frame IDs, the demodulation of the QPSK band was completed. With all the frequency, phase, and temporal information obtained, the Gaussian quantum symbols could also be optimally extracted.
Figure 3 shows various results from experiments conducted with L1–L3 forming a dominant part of the quantum channel [^1] The reasons for excluding results from L4 are explained towards the end of this section..
Figure 3: (Color online) Results from the parameter estimation and calibration procedures performed in the tests carried out at links L1, L2 and L3, where a positive SKR was eventually obtained (refer Table 1 for more info). (a) Number of quantum states successfully measured by RX. (b) Strength of the pilot tone which aids in the clock and carrier recovery. (c) Outcome of the calibration procedure performed after every round of modulation and acquisition. (d) and (e) Noise and loss attributed to Eve. (f) Noise assumed to be beyond Eve’s control. (g) Projected secret key rate (SKR) if the DSP and information reconciliation had been performed in real time. A reconciliation efficiency $\beta=95\%$ and FER = 0.5 was uniformly assumed in the calculations, and for L1 and L2, it was actually also achieved (further details in the main text). The PE was actually performed using all the symbols of Alice and Bob.
@@ 3.2 Protocol runs: Quantum layer
3.3 Protocol runs: Classical layer
@@
With the quantum stage of the protocol finished, the classical data processing steps of IR, PE, and PA were also performed offline. The IR was based on a multi-dimensional reverse reconciliation scheme and a bank of multi-edge-type low-density-parity-check codes were used [34]. For IR, Alice filtered her entire symbol train based on the captured frame ID information, shared by Bob via the classical channel. Due to a limited number of code rates, high-performance IR was actually possible only on data from L1 and L2. In these cases, we successfully obtained a reconciliation efficiency $\beta\approx 95\%$ at a frame error rate (FER) in the vicinity of 0.5.
@@
Alice then performed PE using her filtered symbols and the quantum symbols of Bob recovered by her through IR decoding, together with the variables such as the electronic and vacuum noise variances from the calibrations in RX. After calculating the covariance matrix, the total noise $\xi$ and transmittance $\eta$ were evaluated. These could be partitioned into trusted and untrusted components [15] using the vacuum noise normalization values shown in Fig. 3(c), and the optical transmittance of the receiver (including CWDM-RX) $\eta_{t}=0.33$. Figures 3(d) and (e) show the untrusted components $\eta_{u}=\eta/\eta_{t}$ of the transmittance (converted to loss in dB) and noise $\xi_{u}=\xi-\xi_{t}$ respectively, from the 6 rounds, while the trusted noise values $\xi_{t}$ themselves are shown in Fig. 3(f). Note that these noise figures are as measured, i.e., referred to the output. As is often done in CVQKD, the noise is conveyed as whatever is in excess of the shot/vacuum noise in a relative sense, i.e., in shot noise unit (SNU). Together, these quantities aid in the calculation of the SKR, given by
$\text{SKR}=B(1-\text{FER})\left(\beta I_{AB}-\chi_{BE}\right),$ (1)
where the mutual information $I_{AB}$ between Alice and Bob is also calculated using the covariance matrix. The level of Eve’s correlations to Bob’s data is typified by $\chi_{BE}$, the Holevo information. From a security perspective, the keys generated based on this formula are secure in the asymptotic regime, and assume collective attacks by Eve [35, 23].
Figure 3(g) shows the projected SKR values for L1-L3, which vary mainly due to the fluctuations of the channel parameters $\eta_{u}$ and $\xi_{u}$ during the 6 rounds, and in some cases, also result in the SKR crossing the null threshold ($\beta I_{AB}=\chi_{BE}$). In fact, in case of L4, the values of $\eta_{u}$ and $\xi_{u}$ forced $\beta I_{AB}<\chi_{BE}$ for all rounds, due to which the results from L4 are not plotted in Figure 3. Finally, also note that due to statistical uncertainties $\xi_{u}<0$ in Fig. 3(d) in a couple of cases; for the SKR evaluation, we replaced this with a slightly positive value of 2 mSNU.
4 Discussion
Collecting statistics over the 6 rounds depicted in Fig. 3(g) yields the SKR values as $148.6\pm 94.3$, $434.8\pm 162.8$, and $78.3\pm 81.1$ kbps for L1, L2, and L3, respectively. In this section, we discuss the reasons behind these large variations in the SKR as well as how they can be brought under control. Furthermore, we also provide justification for the assumptions that the DSP, IR, and PA is performed in real time and that it is always possible to obtain $\beta\gtrsim 95\%$ and FER $\lesssim$ 0.5 in IR. With these assumptions, such rates can indeed be achieved in practice.
We also remark that a stronger notion of composability [15] of the secret key could not be satisfied here due to the insufficient number of quantum states obtained in the experiments here as well as the relatively worse trusted loss ($\eta_{t}=0.33$ for qTReX versus $\eta_{t}=0.68$ in Ref. [15]).
4.1 Channel loss: Untrusted vs Physical
@@
The channel loss values graphed in Fig. 3(e) over the 6 rounds show more fluctuation for links L2 and L3 than L1. Furthermore, comparing with the physical loss values presented in Table 1, a difference of at least 0.30.30.3\,0.3dB is evident. The latter can partly be explained due to the fixed loss coming from two extra fiber adapters that were used for connecting to the loopbacks at Odense; since they were absent during the OTDR measurements, their combined loss is not included in the entries of Table 1.
@@
The remaining deviation as well as the fluctuation stems primarily from the subpar performance of the SOP optimization. This is discussed in further detail in subsection 4.1.1 below. From the perspective of channel loss estimation, a poor SOP optimization reduces the efficiency of mode-matching the LLO with the optical signal (containing both the quantum signal and the ancillary signals; see subsection 2.6). This can make the execution of the overall DSP chain [17] at RX inefficient as well: Below a certain pilot SNR, a cascading decrease in performance through the chain is obtained, which worsens the carrier and clock recovery, causing the PE evaluated noise and/or loss to rise.
4.1.1 SNR variations
Figure 4 gives a deeper insight into the varying characteristics of the links (L1-L4) investigated in this work by comparing them with similar results from a 10 km long lab fiber spool (L0), investigated before in Ref. [16].
Figure 4: (Color online) Fluctuations of the relative SNR estimated from the QPSK signal for quantum channels L0-L4. The relative SNR, denoted by $\text{SNR}_{\text{lin}}$ / $\langle\text{SNR}_{\text{lin}}\rangle$, is calculated as the ratio of linearized SNR values from a ($\leq 10~{}$mins long) round to the mean value computed over that entire round, and allows an unbiased comparison among all the channels, where the operating SNRs were otherwise widely different. (a)-(e) Relative SNRs as a function of the IDs of the successfully captured frames from one exemplary round. Each data point in these 5 subplots reflects the SNR calculated from a total of $20$K QPSK symbols, constituting one frame. The frame ID sequence differs in each subplot because of the usage of adaptive trigger level, explained in Section 3.2, which causes the successful-trigger-instants to differ in every round (and thus, obviously also across different channels). (f)-(j) Respective histograms drawn using the relative SNRs from all 6 rounds; the distributions thus reflect the relative SNR variations over a period of 1 hour. The bin width is uniformly chosen to be 0.01 for all links.
More specifically, Figs. 4(a)-4(e) show the variation of the QPSK signal SNR from a single round of the QKD protocol, executed right after the dynamic SOP optimization procedure explained in Section 3.1. The corresponding normalized histograms in Figs. 4(f)-4(j) for the channels are plotted using data accumulated from all 6 rounds, with the number of elements for the histograms, i.e., frames, mentioned in the respective legends.
As outlined in Section 3.2, a block of 20 frames is periodically modulated by the transmitter: Due to this, chunks with up to 20 frames/block may be seen in Figs. 4(a)-4(e). An inter-block variation of the SNR, which occurs on a few $100$s of millisecond scales, is much more pronounced for L2-L4 than either L0 or L1. A vertical elongation of these chunks, indicating much faster fluctuations (on a $1-10\,$ms scale), is most prominent for the longest channel L4. These could be explained by the fact that the rate of change of the SOP of a lightwave in OPGWs can be much higher than in underground fibers, as the former are exposed to not only larger and faster temperature variations but also the natural forces of wind and lightning [36].
Any intra and inter-block variations, together with variations from one round to another, result in the distributions for L2, L3 and L4 being significantly wider than for L0 and L1, as illustrated in Figs. 4(f)-4(j). We conjecture that the different physical locations (Odense vs Lyngby; see Table 1) of the QKD devices play a role in the broadening of the distributions. To elaborate, the QKD devices, specifically the electronic interface of the polarization controller inside CVQKD-TX (see subsection 3.1), were subjected to a high degree of electromagnetic interference (EMI) emanating from the power transformers and grid-lines in case of L2-L4222We note here that measurement results (not shown in Fig. 4) from a B2B configuration at Odense also indicate a wider distribution, however the collected data was an order of magnitude lower, which hinders a reliable statistical conclusion.. In contrast, the experiments conducted over L0 and L1 had the devices placed in a lab environment with a significantly lower EMI.
This implies that the SPSA method used for optimizing polarization works fairly well for links such as L0 and L1, but to actively compensate for fluctuations in L2-L4, the SOP optimization should be run more frequently. Also, with a longer and more thorough characterization of the aerial fibers, a better choice of the initial optimization parameters may help; however, this was not possible due to time constraints. In a future work, it would also be interesting to accurately identify the individual contributions of the SOP fluctuations and high-noise environment to these SNR variations. Finally, an alternative that does not require the implementation of any active method for SOP optimization is to use a so-called polarization hybrid [37, 38], where the fluctuations can be taken into account during DSP.
4.2 Practical and real-time data processing